OpenStudy (anonymous):

SHOW ME HOW TO SOLVE 1/8= X - 7/12

OpenStudy (anonymous):

x=1/8+7/12=17/24 ,BTW,why did you ask such AMAZING question,LUC?

OpenStudy (anonymous):

CAN YU SHOW THE BREAK DOWN IN HOW YOU SOLVE THIS EQUATION

OpenStudy (anonymous):

em,1/8=3/24,7/12=14/24,then X=3/24+14/24=(3+14)/24=17/24,is that clear?

OpenStudy (anonymous):

First of all, you need to combine like terms. This means you need to add 7/12 to both sides giving you 7/12 + 1/8 = x. Your next step is to add the 2 fractions, but of course you need to find the least common denominator, which, in this case is 24. This means that an equivalent fraction for 7/12 would be 14/24 and for 1/8 it would be 3/24. Now to add fractions with common denominators, you add the numerators, 14 + 3, which equals 17. So your answer is 17/24 =x, or, better written, x=17/24.

OpenStudy (anonymous):

THE COMMON DENOMINATORS, ON BOTH GOT IT! THANKS GUYS

OpenStudy (anonymous):

Actually Peachy's explanation is much detailed than me =▽=I'm impressed,really.